OpenStudy (anonymous):

if the measure of each side of square is doubled. its area increases how many times?

OpenStudy (aripotta):

four times

OpenStudy (anonymous):

then if it is tripled: 6 times? @AriPotta

OpenStudy (aripotta):

9 times

OpenStudy (anonymous):

If a square has a side length of 4 centimetres, then its area is equal to 4 x 4 = 16cm2 (16 square centimetres). If a square has a side length of 8 centimetres, then its area is equal to 8 x 8 = 64cm2 (64 square centimetres). Therefore, by doubling the side length of a square, the squares area quadruples.

OpenStudy (anonymous):

4

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

if it is tripled. the solution is 3 x3?

OpenStudy (aripotta):

you square the amount you changed the dimensions. for example, you double the dimensions, so 2 squared is 4. the area is 4 times the original. if you triple the dimensions, 3 squared is 9, so the area is 9 times the original.

OpenStudy (aripotta):

you see what i mean?

OpenStudy (anonymous):

YEA! thanks alot!

OpenStudy (aripotta):

:D

OpenStudy (anonymous):

harry potta? who are you pm me?

OpenStudy (aripotta):

hm?

OpenStudy (anonymous):

i thought you are my classmate xD

OpenStudy (aripotta):

oh lol my name is arielle. everyone calls me ari and i really like harry potter, so i named myself ari potta (like a british accent) :l

OpenStudy (klimenkov):

Let the side of square be \(a\). So, its area will be \(S=a^2\). Now increase this side by \(k\) times and we will get the new side which length is \(k\cdot a\). Find an area of this square with a new side: \(S'=k^2\cdot a^2\). Now find the ratio: \(\frac{S'}{S}=\frac{k^2\cdot a^2}{a^2}=k^2\). To find your answer let k=2.

OpenStudy (anonymous):

HAHAHAHA .. ok xD